{0.2 0.3 0.2

0.4 0.1 0.2

0.1 0.3 0.2}

ย Give three industries having input coefficient matrix as above. Find the production of output of X1 X2 X3 IF THE final demand 10 5 6. Provide economic meaning of 0.1 and 0.4 if any


1
Expert's answer
2021-04-21T07:29:50-0400
"solution"

given the input as :


"\begin{bmatrix}
 0.2& 0.3 & 0.2 \\0.4& 0.1& 0.2\\
 0.1& 0.3& 0.2 
\end{bmatrix}"


production of output"\ X1,X2,X3." is:


"\begin{bmatrix}
 0.2& 0.3 & 0.2 \\0.4& 0.1& 0.2\\
 0.1& 0.3& 0.2 
\end{bmatrix}
\begin{bmatrix}
 X_1 \\
 X_2\\
X_3
\end{bmatrix}=\begin{bmatrix}
 10 \\5\\
 6
\end{bmatrix}"


"X_1=40\\
X_2=65\\
X_3=87.5"


"0.1" or "\frac{1}{10}" , "0.4" or "\frac{4}{10}" are used to show a fraction of a part/ production line used in obtaining the final product / output in this case.